NUCLEAR HYPERFINE STRUCTURE IN THE X^{3}Σ^{+} STATE OF ^{91}ZrC

\underline{S.J. RIXON}, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, B.C. V6T 1Z1, Canada; P.K. CHOWDHURY, and A.J. MERER, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B.C. V6T 1Z1, Canada.

Electronic bands of zirconium monocarbide, ZrC, can be observed following the reaction of laser-ablated Zr atoms with methane under supersonic free-jet conditions. In our experiments some of the bands near 17000 cm^{−1} are strong enough for nuclear hyperfine structure from the ^{91}Zr isotope (I = 5/2, 11.22% abundance) to be assignable. Hyperfine splittings of up to 0.2 cm^{−1} are found in some of the rotational lines. Analysis shows that the principal hyperfine effects are in the ^{3}Σ ground state, where b = −0.03132 ± 0.00015 cm^{−1} and c = −0.00122 ± 0.00038 cm^{−1} (3σ error limits). The large Fermi contact parameter, b, indicates that an unpaired Zr 5sσ electron is present, which, taken together with the small value of λ (0.5139 cm^{−1}), means that the ground state must be a ^{3}Σ^{+} state, from the electron configuration (Zr 5sσ)^{1} (C 2pσ)^{1}. Internal hyperfine perturbations occur between the F_{1} and F_{3} electron spin components of the ground state in the range N = 2 − 4, producing extra lines in some of the branches; the perturbations are of the type ΔN = 0, ΔJ = ± 2, and are a second order effect arising because the F_{1} and F_{3} spin components (J = N + 1 and J = N − 1, respectively) both interact with the F_{2} component (J = N) through ΔN = 0, ΔJ = ± 1 matrix elements of the Fermi contact operator.